OpenStudy (anonymous):

How many solutions are there to the following system of equations? 5x + 3y = 7 2x – y = 2 A. infinitely many B. 1 C. 0 D. 2

OpenStudy (anonymous):

Try solving the those equations either by substitution method or elimination method then see how many solutions u get

OpenStudy (amoodarya):

2x – y = 2 2x-2=y so put y=2x-2 in first equation to solve for x 5x + 3y = 7--->5x + 3(2x-2) = 7